# Elastic Sketch under Random Stationary Streams: Limiting Behavior and Near-Optimal Configuration

### 17 Mar 2026 v1

@@ 2 Problem Formulation
The update of bucket $b(t)$ depends on these three cases. In case 1, $S_{b(t)}$ starts/keeps monitoring $R_{t}$, and $V_{b(t)}^{+}$ is incremented by $1$. In case 2, the bucket $b(t)$ keeps monitoring $S_{b(t)}(t-1)$, and in this case it is $V_{b(t)}^{-}$ that is incremented by $1$ and not $V_{b(t)}^{+}$. In case 3, $S_{b(t)}$ changes value to $R_{t}$, and the counts in $V_{b(t)}^{+}$ and $V_{b(t)}^{-}$ are reset to $1$ and 0, respectively. Note that by construction of the algorithm, it is always true that $\lambda V_{b}^{+}(t)-V_{b}^{-}(t)\geq 0$, and thus all cases are covered. Formally,
@@
On the other hand, when $m_{1}>0$, the update of $\bm{Y}$ depends on the three cases distinguished for the heavy block. The matrix $\bm{Y}$ is not updated on hits to the currently elected item at bucket $b(t)$ (case 1), but it is updated on arrivals of items mapped to $b(t)$ that are not $S_{b(t)}(t-1)$. In case 2, $R_{t}$ is added to the sketch with value $1$. Upon an eviction of the elected item at bucket $b(t)$ (case 3), the epoch ends: $V_{b}^{+}(t-1)$ holds the evicted item’s total during the epoch, so this amount is added to the sketch. Formally, for any counter $(l,c)$,
@@ 5.1 Definitions and Notation
In words, upon a request for an item not in the bucket in question, $M_{b}(t)$ remains unchanged (w.p. $1-\mu_{b}$). Moreover, when $U_{b}(t)$ is positive, $S_{b}(t)$ remains unchanged, and either $U_{b}(t)$ increases by $\lambda$ (with probability $p_{S_{b}(t)}$) or decreases by $1$ (w.p. $\mu_{b}-p_{S_{b}(t)}$). Once $U_{b}(t)$ hits 0, $S_{b}(t)$ may switch to a new item $i$ w.p. $p_{i}$ and $U_{b}$ increases by $\lambda$ anyway. Figure 3 illustrates the one step transitions of this Markov process when the total number of items is $2$ and the number of buckets is $1$. We denote the set of states with $U_{b}=0$ as $\overline{\mathbf{0}_{b}}$, i.e., $\overline{\mathbf{0}_{b}}\triangleq\{(i,0):\;i\in\mathcal{I}_{b}\cup\{-1\}\}$.
@@
where $\phi(x,\lambda)=\sum_{k=0}^{\lambda}x^{k}$. For $x\in[0,1]$, $\phi(x,\lambda)$ takes values in $0,\lambda+1]$, and it is strictly increasing in $x$. Thus, (i) when $\lambda\leq\frac{\mu}{p}-1$, $1$ is the unique root of $f$, and therefore $r_{n}=1$, $\forall n$, i.e., $G$ is recurrent, and (ii), when $\lambda>\frac{\mu}{p}-1$, $f$ has two distinct roots, $1$, and $x_{0}\in(0,1)$, such that $\phi(x_{0},\lambda)=\mu/p$. Next, we prove that $r_{1}<1$ when $\frac{\mu}{p}<\lambda+1$, to deduce that $r_{1}$ is the solution to the equation $\phi(\cdot,\lambda)=\mu/p$.
@@
When $n_{b}=1$, $M_{b}$ is transient, because $U_{b}$ diverges w.p. $1$.
@@
where we have used the fact that $\lim_{t\to\infty}N_{i}(t)/t=p_{i}$, and $N_{i}(T_{b}^{\text{evict}}(t))\leq T_{b}^{\text{evict}}$ for any item $i$. On the other hand, for any $b^{-}\in\mathcal{B}_{\beta}^{-}(\lambda)$, we have that a.s.,
@@ 6 Conclusion
where $\delta_{\ell,c}(i)=\mathds{1}(h_{l}(i)=c)$. Since $Y_{\ell,c}^{\texttt{CM}}(t)=\sum_{i\in\mathcal{I}}\delta_{\ell,c}(i)N_{i}(t)$, the difference $Y_{\ell,c}^{\texttt{CM}}(t)-Y_{\ell,c}(t)$ from (69) can be partitioned into $m_{1}$ terms each corresponding to items in $\mathcal{I}_{b}$. In each one of these partitions, $\delta_{i}(t,\beta)$ is only equal to $1$ for $S_{b}(t)$ and hence (66) is equivalent to (69). The proof of (69) follows an induction argument.
@@
Case 3: $S_{b}(t-1)\neq R_{t}$ and $\lambda V_{b}^{+}(t-1)-V_{b}^{-}(t-1)>0$ (mismatch, vote passes).  The bucket content is not modified ($S_{b}(t-1)=S_{b}(t)$, $T_{b}^{\text{evict}}(t-1)=T_{b}^{\text{evict}}(t)$, and $V_{b}^{+}(t)=V_{b}^{+}(t-1)$), only $V_{b}^{-}$ increases, and $\bm{Y}$ is updated by incrementing the cells corresponding to $R_{t}$ by $1$. Thus both sides of equation (65) remains unchanged. Moreover, $\delta_{R_{t}}(t,\beta)=\delta_{R_{t}}(t-1,\beta)=0$ and hence the RHS of (66) for all cells $(\ell,h_{\ell}(R_{t}))$ with $r\in[d]$ increase by $1$ and thus (66) holds at step $t$.
@@
which enables us to deduce (85). Note that $x\mapsto\lambda x^{\lambda+1}-(\lambda+1)x^{\lambda}+1$ is decreasing over (0,1)(0,1)—direct computation of the derivative—and its value at 0 and $1$ are positive, and thus $\partial_{\lambda}r\leq 0$ and $\partial_{C}r\geq 0$.

### 26 Mar 2026 v2

@@ 2 Problem Formulation
The update of bucket $b(t)$ depends on these three cases. In case 1, $S_{b(t)}$ starts/keeps monitoring $R_{t}$, and $V_{b(t)}^{+}$ is incremented by 11. In case 2, the bucket $b(t)$ keeps monitoring $S_{b(t)}(t-1)$, and in this case it is $V_{b(t)}^{-}$ that is incremented by 11 and not $V_{b(t)}^{+}$. In case 3, $S_{b(t)}$ changes value to $R_{t}$, and the counts in $V_{b(t)}^{+}$ and $V_{b(t)}^{-}$ are reset to 11 and 0, respectively. Note that by construction of the algorithm, it is always true that $\lambda V_{b}^{+}(t)-V_{b}^{-}(t)\geq 0$, and thus all cases are covered. Formally,
@@
On the other hand, when $m_{1}>0$, the update of $\bm{Y}$ depends on the three cases distinguished for the heavy block. The matrix $\bm{Y}$ is not updated on hits to the currently elected item at bucket $b(t)$ (case 1), but it is updated on arrivals of items mapped to $b(t)$ that are not $S_{b(t)}(t-1)$. In case 2, $R_{t}$ is added to the sketch with value 11. Upon an eviction of the elected item at bucket $b(t)$ (case 3), the epoch ends: $V_{b}^{+}(t-1)$ holds the evicted item’s total during the epoch, so this amount is added to the sketch. Formally, for any counter $(l,c)$,
@@ 5.1 Definitions and Notation
In words, upon a request for an item not in the bucket in question, $M_{b}(t)$ remains unchanged (w.p. $1-\mu_{b}$). Moreover, when $U_{b}(t)$ is positive, $S_{b}(t)$ remains unchanged, and either $U_{b}(t)$ increases by $\lambda$ (with probability $p_{S_{b}(t)}$) or decreases by 11 (w.p. $\mu_{b}-p_{S_{b}(t)}$). Once $U_{b}(t)$ hits 0, $S_{b}(t)$ may switch to a new item $i$ w.p. $p_{i}$ and $U_{b}$ increases by $\lambda$ anyway. Figure 3 illustrates the one step transitions of this Markov process when the total number of items is 22 and the number of buckets is 11. We denote the set of states with $U_{b}=0$ as $\overline{\mathbf{0}_{b}}$, i.e., $\overline{\mathbf{0}_{b}}\triangleq\{(i,0):\;i\in\mathcal{I}_{b}\cup\{-1\}\}$.
@@
where $\phi(x,\lambda)=\sum_{k=0}^{\lambda}x^{k}$. For $x\in[0,1]$, $\phi(x,\lambda)$ takes values in $0,\lambda+1]$, and it is strictly increasing in $x$. Thus, (i) when $\lambda\leq\frac{\mu}{p}-1$, 11 is the unique root of $f$, and therefore $r_{n}=1$, $\forall n$, i.e., $G$ is recurrent, and (ii), when $\lambda>\frac{\mu}{p}-1$, $f$ has two distinct roots, 11, and $x_{0}\in(0,1)$, such that $\phi(x_{0},\lambda)=\mu/p$. Next, we prove that $r_{1}<1$ when $\frac{\mu}{p}<\lambda+1$, to deduce that $r_{1}$ is the solution to the equation $\phi(\cdot,\lambda)=\mu/p$.
@@
When $n_{b}=1$, $M_{b}$ is transient, because $U_{b}$ diverges w.p. 11.
@@
where we have used the fact that $\lim_{t\to\infty}N_{i}(t)/t=p_{i}$, and $N_{i}(T_{b}^{\text{evict}}(t))\leq T_{b}^{\text{evict}}$ for any item ii. On the other hand, for any $b^{-}\in\mathcal{B}_{\beta}^{-}(\lambda)$, we have that a.s.,
@@ 6 Conclusion
where $\delta_{\ell,c}(i)=\mathds{1}(h_{l}(i)=c)$. Since $Y_{\ell,c}^{\texttt{CM}}(t)=\sum_{i\in\mathcal{I}}\delta_{\ell,c}(i)N_{i}(t)$, the difference $Y_{\ell,c}^{\texttt{CM}}(t)-Y_{\ell,c}(t)$ from (69) can be partitioned into $m_{1}$ terms each corresponding to items in $\mathcal{I}_{b}$. In each one of these partitions, $\delta_{i}(t,\beta)$ is only equal to 11 for $S_{b}(t)$ and hence (66) is equivalent to (69). The proof of (69) follows an induction argument.
@@
Case 3: $S_{b}(t-1)\neq R_{t}$ and $\lambda V_{b}^{+}(t-1)-V_{b}^{-}(t-1)>0$ (mismatch, vote passes).  The bucket content is not modified ($S_{b}(t-1)=S_{b}(t)$, $T_{b}^{\text{evict}}(t-1)=T_{b}^{\text{evict}}(t)$, and $V_{b}^{+}(t)=V_{b}^{+}(t-1)$), only $V_{b}^{-}$ increases, and $\bm{Y}$ is updated by incrementing the cells corresponding to $R_{t}$ by 11. Thus both sides of equation (65) remains unchanged. Moreover, $\delta_{R_{t}}(t,\beta)=\delta_{R_{t}}(t-1,\beta)=0$ and hence the RHS of (66) for all cells $(\ell,h_{\ell}(R_{t}))$ with $r\in[d]$ increase by 11 and thus (66) holds at step $t$.
@@
which enables us to deduce (85). Note that $x\mapsto\lambda x^{\lambda+1}-(\lambda+1)x^{\lambda}+1$ is decreasing over (0,1)(0,1)—direct computation of the derivative—and its value at 0 and 11 are positive, and thus $\partial_{\lambda}r\leq 0$ and $\partial_{C}r\geq 0$.